OpenStudy (anonymous):

identify the axis of symmetry for -x^2+4x

OpenStudy (anonymous):

Let y=-x^2+4x=-(x^2+4x+4-4)=-(x^2+4x+4)+4=-(x+2)^2+4 axis of symmetry is x+2=0 or x=-2